OpenStudy (anonymous):

Differentiate g(u)=sqrt(2)u+sqrt(3u) is the answer... g'(u)=(sqrt(3)/2sqrt(u))+sqrt(2)

OpenStudy (shaik0124):

y = sqrt(x) use the power rule: deravite of : x^(n) is : n*x^(n - 1) Rewrite: y = sqrt(x) as : y = x^(1/2) y' = (1/2)(x^(.5 - 1)) = (1/2)(x^(-1/2)) y' = 1/(2*sqrt(x) ) follow same procedure u will get answet

OpenStudy (ghazi):

\[\sqrt[2]{U}+ \sqrt {3U}\] ??

OpenStudy (anonymous):

yes.

OpenStudy (shaik0124):

no

OpenStudy (shaik0124):

i think it is 1/2root u+1/root3u

OpenStudy (anonymous):

am I wrong..?

OpenStudy (ghazi):

but monore has said that

OpenStudy (anonymous):

I got... sqrt(3)/2sqrt(u)+sqrt(2)

OpenStudy (anonymous):

(sqrt(3)/2sqrt(u))+sqrt(2)

OpenStudy (anonymous):

@jim_thompson5910 am I correct?

OpenStudy (ghazi):

\[\frac{ d (u)^{5/2}}{ du }+\frac{ d(\sqrt{3u}) }{ du }\]

OpenStudy (ghazi):

@monroe17 after differentiating the above equation you'll get the required answer

OpenStudy (anonymous):

sqrt(3)/2sqrt(u)?

OpenStudy (shaik0124):

answer is differentiation of root xis 1/2root x as i expalined earlier through power rule and differentiation of root3x is 1/6root3u finally 1/2rootx+1/6root3u

OpenStudy (anonymous):

I'm so confused :(

OpenStudy (shaik0124):